### v1

# Disentangling Brillouin’s negentropy law of information and Landauer’s law on data erasure

Abstract
The link between information and energy introduces the observer and his knowledge into the understanding of a fundamental quantity of physics. Two approaches compete to account for this link, Brillouin’s negentropy law of information and Landauer’s law on data erasure, which are often confused. The first, based on the Clausius’ inequality and Shannon’s mathematical results is very robust, while the second, based on the simple idea that information needs a material embodiment (data-bits) is today perceived as more physical and prevails. In this paper, we show that Landauer’s idea results from a confusion between information (a global emergent concept) and data (a local material object). This confusion leads to many inconsistencies and is incompatible with thermodynamics and information theory. The reason it prevails is interpreted to be due to a frequent tendency of materialism towards reductionism, neglecting emergence and seeking to eliminate the role of the observer. A paradoxical trend given that it is often accompanied by the materialist idea that all scientific knowledge nevertheless originates from observation. Information and entropy are actually emergent quantities introduced in the theory by convention.
I Introduction
The definition of something is a statement that allows us to recognize it when we see it. For example, it can be the statement of all the characteristics of that thing. Or, it can be the statement that this thing is the name given to a category of elements already defined. Of course, this supposes that the list of all elements in this category is known, otherwise we would be unable to recognize every time an element as belonging to this category.
For energy, there is no such definition, neither of the first type nor of the second. This is why “in physics today, we have no knowledge of what energy is” (R. Feynmann [1]). As a working definition, one can say that “in the everyday world outside of the formal language of physics, energy is the ability to do anything… Insofar as “does” means “produces a change in what was before”, energy is philosophically the determinant of all observable change”(E. Hecht) [2]). With this qualitative definition, there is no doubt that knowledge and information would be a priori forms of energy. An example is given by Maxwell [3]: an intelligent being (a demon) produces mechanical work from the information they get about a device. If energy cannot come from nowhere according to the principle of conservation, we see that, a posteriori as well, information is energy.
How to be quantitative about information and its conversion into unit of energy? The solution was given by Shannon’s information theory [4] which, by the intermediate of Gibbs statistical mechanics [5], allowed us to identify the entropy of a system as the uncertainty we have about it. Hence the Brillouin’s law of information [6, 7]: decreasing the uncertainty by acquiring information has a minimum energy cost equal to that of the corresponding decrease in entropy.
Is that enough? No, according to Landauer. Energy is a physical quantity, information must be too [8, 9, 10]. But Landauer means by physical, something that is not abstract but tangible and material, in other word an hardware set of data-bit. Being materialized, information (understand data-bits) must be erased prior to acquisition. Hence, the Landauer’s idea: the energy cost for the acquisition of information is paid by this preliminary erasure step. This is the Landauer’s law on data-erasure.
Here, we will see that what may appear as an addition or improvement to the triplet (thermodynamics-statistical mechanics-information theory) is in reality totally incompatible with it. The abstract concept of information cannot be confused with its material support (data-bit), without introducing many inconsistencies into this theory.
The article is organized as follows. The first part provides a brief overview of this threefold theory (thermostatistics) which in fact forms a whole. Particular attention will be paid to the role assigned to the observer and to conventions. Next, the link between information and energy as established by the demonic engines is presented, with Brillouin and Landauer’s views on how they work. Both are often confused (e.g.[11, 12]), so that this part aims in particular to clarify their differences.
Energy is a concept related to the physical world, while information refers to us as intelligent beings endowed with a mind. Even materialists (like myself) have to force themselves to connect the two. That is why, in my opinion, it is an illusion to think that the different conceptions of what science is, of what a theory is, play no role in this problem. Even though physicists often wish to distance themselves from philosophy, an epistemological approach is necessary to understand the interplay between information, energy, probabilities and the role of the observer. Epistemological positions are always present in the literature on this subject, and in particular in that of Landauer. In most cases, they are hidden or implicit. This does not favor clarity.
We also adopt an epistemological position throughout this article, but it will be explicit from the outset. It will be that of neo-positivists [13, 14, 15, 16] and those who influenced them [17, 18, 19, 20]. In brief: 1) only two sources of knowledge are admitted, logic and observation; 2) an explanation is a deduction of observations (past and future) from theory; 3) to fill the gap in our ignorance, the latter must include conventions chosen according to their convenience and to the economy of thought they allow. The last part of the paper shows how this position permits to analyze Landauer’s reductionist idea (beyond its inconsistencies) and how to include emergentism into materialism.
II Thermostatistics
II.1 Thermodynamics: principle versus law
Energy is one of the few concepts in physics with universal scope. So that every scientist knows the fundamental statements on which the theory of thermodynamics, the science of energy transformations, is based. In short:
Definition: The state of equilibrium (or state, or equilibrium) of a system is the stationary situation where no change occurs.
1st principle: The total quantity of energy of an isolated system (the universe) is conserved during any change.
2nd law: There exists a state quantity, which we call entropy $S$, such that for any change from state A to state B, $\Delta S_{\text{AB}}=(S_{\text{B}}-S_{\text{A}})\geq\int\mathop{}\!\mathrm{d}Q/T$ (Clausius inequality), where $Q$ and $T$ are the heat exchanged and temperature, respectively.
But what people are less aware of is the fundamental difference in nature between the last two statements, hence the terms “principle” and “law” used to highlight this point (this denomination follows that of Poincaré [18]). A principle is a convention, a law is inferred from induction. Let us clarify this point starting with the second law.
Clausius inequality expresses that, for a given change of state along a differentiable path, the heat (in temperature unit) received by the system from the environment never exceeds a certain limit that depends solely on the initial and final states. This limit is the difference in entropy. It is usually approached as the rate of change along the trajectory is slowed down in a quasi-static manner. The paths that allow this are called reversible because they can be followed in both directions. The others, especially those which are not differentiable, are called irreversible. In Clausius inequality, the sign of $Q$ refers to the system (positive when received). For simplicity, consider a system with constant internal energy, typically a set of independent entities, such as an ideal gas or a set of bits at a given temperature. Then for any change, transfers of mechanical work ($W$) and heat ($Q$) with the environment compensate each other ($\mathop{}\!\mathrm{d}Q=-\mathop{}\!\mathrm{d}W$), so that the Clausius inequality rewrites by considering the reverse change from B to A: $\Delta S_{\text{BA}}\leq\int\mathop{}\!\mathrm{d}W/T$. It follows that the entropy difference essentially corresponds to the observed minimum work (in temperature unit) required for the system to return to its initial state by a quasi-static path (hence the word “entropy” from Greek entropia “a turning toward”). It is clear that mechanical work and heat can be measured in as many different circumstances as one can imagine. The Clausius inequality results from the generalization of these specific observations, the regularity of which has never been contradicted in two centuries of experiments in this area. The case of equality allows us to define the word “entropy”, but this is just a nominal definition, the core of the second law lies in the inequality. The second law follows from induction.
As for the first principle, it states that the energy of the universe is conserved regardless of the change undergone by a system. In other words, the energy balance is always zero. One could imagine that the first principle also derives from precise measurements of the energy balance of various changes, which would lead to the conclusion that it is apparently always zero, thus allowing these observations to be generalized and elevated to the rank of a universal law. Of course, this assumes that we are able to recognize (identify) energy, whatever form it takes, when we see it. But that is not how it happened, because there is absolutely nothing that allows us to recognize something a priori as being a form of energy before this form has been added to the list of those known. And this is precisely this kind of measurements that allows for such an addition. The first principle does not follow from inductive reasoning. It is actually the hidden definition of what energy is, because there is no other [1].
Although this would be extremely unlikely for the second law, there is no conceptual impediment to discovering a counterexample that would invalidate the generality of any inductive law. But for a definition, such as the first law, such an invalidation is conceptually meaningless. A definition cannot be invalidated by an observation: “[A] principle…is no longer subject to the test of experiment. It is not true or false, it is convenient”[19]. In case an experiment would lead to a non-zero energy balance, we have two options: 1) either we decide that the 1st principle is no longer appropriate; 2) or we acknowledge that something escaped to us and that we have just discovered a new form of energy. Clearly, the second option is more convenient, and that is how all forms of energy were added to the list of those known [14], for instance kinetic energy (energy of motion) or rest mass (energy at rest). Ultimately, energy is defined only as a category comprising an undetermined number of items, a number that depends on the state of the art. Energy is an abstract conventional concept, it is like a universal in the metaphysical sense given to this word. But this universal has a particularity: it includes a number of elements that depends on our knowledge. So that internal energy cannot be considered as intrinsic to the state of a system, nor as intrinsic to certain phenomena. That is to say, energy cannot be considered independent of the observer. “Energy is […] the determinant of all observable change” (E. Hecht [2]).
What about entropy? The notion of change is central in the first principle defining energy (energy is conserved when everything else change), but also in the second law defining entropy. The latter is not defined in an absolute way by the Clausius’ inequality, but only in a relative way from the measure of the heat exchanged when the system undergoes a change of state. No observable change, no difference of entropy. We could say the same for any state quantity, for example temperature. But for temperature, we have means to measure its value for a given state without the need of any change. For entropy there is no such mean. Entropy is a state quantity that can only be deduced from the observation of a change of state. Ultimately, entropy depends on the ability of the observer to observe these changes, that is to say from the ability of the observer to perceive differences. No work is needed to return to the initial state if no change is observed: “The idea of dissipation of energy depends on the extent of our knowledge” (J. C. Maxwell [21]).
Like internal energy, entropy also depends on the observer’s knowledge and cannot be considered an intrinsic property of the state of a system.
II.2 Statistical mechanics: reductionism versus emergentism
Statistical mechanics, “the rational foundation of thermodynamics”[5], aims to remedy the subjectivity introduced by the observer in thermodynamics. The goal is to calculate everything from the Newton’s mechanics of atoms, that is to say from the behavior of objects totally independent of us. It is a reductionist approach.
At the atomic scale everything is in motion. Microscopic configurations (also named phases or microstates) are constantly changing. But they are also so numerous that it is impossible to hope to know that of a system other than in terms of probabilities. So, statistical mechanics faces two main problems:
  1. 1.
    Prior distribution: What phase probability distribution should be used as a starting point for calculations, given that it cannot be measured?
  2. 2.
    Equilibrium: Since phases are never stationary, a definition of equilibrium other than that of thermodynamics is necessary. Which one?
The first problem is generally solved by the ergodic hypothesis, which essentially considers that the phases distribution of snapshots of identical systems (which differ only in their phase) is equal to that of a single system evolving in time. The trajectory of the system in the phase space $\Gamma$ is supposed to be volume preserving, like that of a deterministic mechanical system (Liouville’s theorem). For two successive points $a$ and $b$ of the trajectory, if the system is deterministic the probability for the system to be in $b$ knowing it was in $a$ is equal to 1. So that the probability for $b$ is equal to that of $a$, and this is true throughout the trajectory. This results in a uniform probability distribution of phases for isolated systems, thus opening the door to additional calculations and great results. In particular the Boltzmann distribution for a closed system and the temperature dependence of the partition function [22], which by identification to known equations of thermodynamics leads finally to the famous Gibbs’ result for entropy:
$S=\sum\limits_{i\in\Gamma}p_{i}\ln 1/p_{i}$ (1)
where $p_{i}$ is the probability of phase $i$. In the case where the distribution is uniform with $p_{i}=1/\mathcal{W}$, one has
$S=\ln\mathcal{W}$ (2)
Clausius entropy, initially defined in thermodynamics solely by its measurable variations, was equal to the minimum amount of heat received by the environment. Thanks to statistical mechanics, it is now understood in absolute terms as the logarithm of the number of possibilities for the microscopic configuration. If understanding means creating connections, then progress is huge.
But getting over the observer is not without introducing new problems. Where has the observer’s knowledge gone, the information he possesses? The best illustration of this problem is given by the famous two Gibbs paradoxes [23, 24]:
  1. 1.
    Joining two identical volumes of the same gas increases the volume accessible to each particle and therefore the total number of possibilities for the system and its Gibbs entropy. However, this occurs without heat exchange and thus without variation of Clausius entropy. The system can return to the initial state at no work, simply by replacing the partition between the two volumes.
  2. 2.
    Mixing two volumes of gas requires work to return to the initial state only if these two gases were initially identified as different. But for statistical mechanics, both cases increases entropy because replacing the partition between the two volumes is not enough to ensure that each particle returns to its original compartment.
Actually in these two problems, statistical mechanics consider the overall information needed to describe the system (which particle in which compartment), whereas thermodynamics consider an incomplete information. “It is to states of systems thus incompletely defined that the problems of thermodynamics relate” (J.W. Gibbs  [25]).
In thermodynamics, the perception of a change depends not only on the observer’s knowledge, but also on what they consider relevant. In a certain sense, it is true that “no man ever steps in the same river twice” (Heraclitus). But on the other hand, what makes the identity of the river and ours is not that of the molecules that constitute us: tomorrow, it will still be the Seine that flows through Paris and I hope to still be myself. The identity of molecules is information that exists, at least for large traceable molecules, but it is considered in thermodynamics by the observer as meaningless or irrelevant in certain cases (e.g. for open systems). The level of information considered relevant is a convention. And it is only in this way, by reintroducing the observer and the information they consider relevant, that Gibbs paradoxes can be resolved for large traceable molecules (see [26] p.13-14 and [27]).
The second problem of statistical mechanics is related to the definition of equilibrium. The current phase is constantly changing. We could simply define the equilibrium as the macroscopic state whose properties take the average values of those of the microscopic configurations. But this is not enough, something is missing. From experience, I know that the equilibrium state of a gas in a room is that in which it uniformly occupies the entire volume and not just a corner. Fluctuations occur which take the system away from equilibrium, but there is something that restores it. Consider a gas inside a cylinder below a piston. We can push or pull the piston and experience a force (Figure 1). The system behaves as a metal spring would. The difference is that for the metal spring, the macroscopic equilibrium reflects the microscopic equilibrium of the atoms that are in a potential well. Such a potential well does not exist for the atoms of an ideal gas. One could argue that the equilibrium position of the piston corresponds to the equality of internal and external pressures, and calculate the pressure as being related to the number of atomic collisions on both side. In doing so, as for the metal spring, macroscopic forces would ultimately be explained by microscopic ones. However, the calculation inevitably starts with the assumption that the atom density near the surface is equal to that of the gas at equilibrium for a given piston position. In other words, the calculation starts with the very assumption it intends to prove. This is circular reasoning.
In reality, the force exerted on the piston is an emergent property, just like pressure and entropy. It cannot be derived consistently from the microscopic scale and the laws we already have at our disposal. An additional ingredient is necessary that must be postulated upstream in the theory by the conventional definition we give of macroscopic equilibrium. Just as a metal spring minimizes its potential energy at equilibrium, it would be convenient if a volume of gas would optimize something. In thermodynamics we already has the second law that says that $\mathop{}\!\mathrm{d}S\geq 0$ for an isolated system. If we postulate that entropy is maximum at equilibrium the trick is done [28] and we have the restoring force we are looking for. The equilibrium would become an attractor.
Figure 1: A piston compressing either a metal spring (left) or a volume of gas (right) experiences a restoring force when it moves away from equilibrium. For the metal spring, this force has a microscopic origin. For the gas, it is emergent.
This is where the problem arises with such a postulate within the framework of statistical mechanics as we have presented it so far. According to Poincaré’s recurrence theorem [29] a volume preserving dynamical system is recurrent and has no attractor. This is inherent to the ergodic hypothesis and its conception of probabilities as frequencies of occurrence: frequency implies recurrence. The solution would consist in getting out the ergodic hypothesis. But then comes back the problem of the prior probability distribution of phases. It is a vicious circle. Finally, the best approach would be to admit that “When one does not know anything the answer is simple. One is satisfied with enumerating the possible events and assigning equal probabilities to them.” (R. Balian [30]). This is known as the fundamental postulate of statistical mechanics but in reality nothing more than the Laplace’s “principle of insufficient reason”[31]. The problem is that, as it stands, it is unfounded. It is a synthetic a priori knowledge: “It cannot be that because we are ignorant of the matter we know something about it” (R.L. Ellis [32]).
Information theory provides us precisely the missing pieces that we lack. It reintroduces the observer’s knowledge (because information is meaningless without an observer) and solves the problem of prior probabilities by making the principle of insufficient reason analytical.
II.3 Information theory: come back of the observer
Shannon [4] tackled a problem that appeared very different from that of thermostatistics, but which ultimately turned out not to be so far: the lossless compression of a message. His started by getting out the message meaning and by considering its emitter as a source of a random variable taking its values among a set $\Gamma$ of possible characters. In doing so, he mathematically demonstrated two results:
  1. 1.
    In no case the average number of bits per character can be less than
    $H=\sum_{i\in\Gamma}{p_{i}\log_{2}(1/p_{i})}$ (3)
    $H$ is named quantity of information emitted by the source and by identification with Eq.1, $S=H\times\ln 2$ is its entropy.
  2. 2.
    Within a factor, $H$ (and thus $S$) is the only measure of uncertainty on the upcoming character that is 1) continuous in $p$; 2) increasing in $\mathcal{W}=1/p$ for uniform distributions; 3) additive over different independent sources of uncertainty.
These results are general and apply regardless of the random variable, and in particular to the constantly evolving phases of any physical dynamic system. It follows that, having been a quantity of heat dissipated (Clausius), then the logarithm of a number of possibilities (Gibbs), entropy is now (Shannon), and all at once, the uncertainty that the observer has concerning the phase of the system. Here again, its understanding had made great progress.
These two results were quickly recognized as a major advance for thermostatics. Brillouin, with his law of information [6, 7] and Jaynes, with the maximum entropy principle [33, 34], were the firsts. The first point will be presented in the following section; we will therefore give a brief overview of the second here.
How to describe or mentally represent, in a rational way, something that we only partially know? That is in reality the central question of science. The description must account for all pieces of information, otherwise it would be incomplete, but it must not invent an information that comes out of nowhere (which would constitute a synthetic a priori knowledge). For instance, this is the problem of linear-regression: for the solution (the description) to be unique, the degree of the polynomial must be less than the number of points. Solving such a problem amounts to seeking a unique description that maximizes the uncertainty about the information we do not have. For a description involving a probability distribution, Shore and Johnson [35] showed that maximizing its entropy (instead of another possible measure of uncertainty) is the only procedure leading to a unique solution.
Hence the theorem of maximum entropy: the best prior probability distribution that accounts for our knowledge is that of maximum entropy. For instance, Shannon [4] showed that “when one does not know anything…” (Balian) the maximum of entropy is obtained for a uniform distribution. This provides a mathematical basis for Laplace’s principle of insufficient reason and makes it analytical.
By getting ride of the ergodic hypothesis, there is absolutely no longer any obstacle, no inconsistency, in postulating by convention that equilibrium is an attractor, that is to say, in defining equilibrium as the state of maximum entropy of phases distribution (that is also the state of maximum entropy of variables whose distributions are similarity-invariant in form [36], such as the density). This is the principle of maximum entropy [33, 34]. This principle, like any other definition, is conventional. It is not required to be checked by experiment. And thankfully so, because it would not be possible. This principle is just convenient. Without introducing inconsistency, it allows us to establish a deductive link between theoretical statements (the definition of equilibrium plus the 2nd law) and experiments (the observed restoring force towards equilibrium), that is to say, to explain and account for the latter.
III Information and demons
III.1 From Maxwell to Szilard
The connection between information and energy was established by Maxwell [3]. He imagines an intelligent being (say a demon or a computer) capable of tracking particles and, from this knowledge, extracting energy from the system which would otherwise be impossible. A simpler version of such an engine is that of Szilard [37], shown in Figure 2. Applying the first law of thermodynamics, which states that energy cannot by definition come from nowhere, forces us either to admit that information is a form of energy or to formulate another definition of energy. Undeniably, the first alternative is more convenient.
However, it is necessary to be more quantitative and explicit. This is precisely the purpose of Brillouin’s law of information and also that of Landauer’s about data erasure.
Figure 2: Cycle of operations of Szilard’s engine. By knowing the location of the particle, a demon (say a computer) can produce mechanical work with no other cost than that of acquiring the information. The first principle implies that information is energy.
III.2 Brillouin’s negentropy law of information
Brillouin’s reasoning consists of three premises:
  1. 1.
    Clausius: The negative of the entropy difference $\Delta S$ experienced by a system (at given $T$) is the minimum work $W$ that must be done on the system for the change to take place: $W\geq-T\Delta S$.
  2. 2.
    Gibbs: Entropy $S$ is related to the probability distribution of possible microstates.
  3. 3.
    Shannon: The Gibbs formula for $S$ is actually to a factor $\ln 2$ that of the uncertainty $H$ on the actual microstate: $S=H\ln 2$
and one deduction:
  1. 4.
    Brillouin: Therefore, reducing uncertainty by acquiring information requires minimum work: $W\geq-T\Delta H\ln 2$. For one bit of acquired information $\Delta H=-1$, so that:
    $W_{\text{acq/bit}}\geq T\ln 2$ (4)
where $W_{\text{acq/bit}}$ is the minimum work that we have to provide (and that will be dissipated as heat) per bit of acquired information. This is the Brillouin’s negentropy law of information [6, 7], called “principle” by Brillouin [6], also called sometime Szilard’s principle [38] (because Eq.4 can be derived from the functioning of Szilard engine), but here called “law” for the sake of consistency with §II.1 (because it is derived from the 2nd law). With Brillouin’s equation, the energy balance of Szilard engine is zero, as required.
Let us consider a body of mass $m$, lifted by a height $\Delta h$. Classical mechanics tell us that the increase in potential energy is $mg\Delta h$. This result do not impose any particular steps for the change in height and do not specify where the mechanical work is done. However, the result concerning the net change in potential energy is general and valid in all cases, regardless of the path taken.
Brillouin’s law works in exactly the same way. Entropy, and consequently the uncertainty about a system or the information we lack to describe it, is a state quantity whose variation is independent of the path connecting two states. Thus, Brillouin’s law does not tell us where the energy cost of acquisition is paid in the operation. And this is inherent to the definition of what is a state quantity. Acquisition may not be a simple operation, but may consist of other more elementary operations, which may differ depending on the specific data acquired. But at the end, whatever the way acquisition is done, it will cost at least $T\ln 2$ in average per bit.
Brillouin’s law is a syllogism and says nothing more than its premises, in this case two mathematical results and the second law. It follows that its generality is as robust as that of the second law.
III.3 Landauer’s law on data erasure
Landauer’s competing argument [8] is based on two main ideas. The first is that “information is not a disembodied abstract entity; it is always tied to a physical representation. It is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent”[10]. In other words, information is always supported by a hardware with two discernible different configurations (two values 0 or 1), namely a data-bit. The second idea is that the acquisition of one bit of information necessarily passes by the erasure of its supporting data-bit. The reasoning is the following:
  1. 1.
    Any intelligent being has a finite memory, so that the infinite cyclic acquisition of information about a dynamical system necessarily requires the erasure of data-bits.
  2. 2.
    Erasing a data-bit (a thermodynamical system) consists of setting it to an arbitrary value (say 0). The procedure must be able to work for a known or unknown initial value (i.e. it must be the same in both cases). This constraint automatically implies a two-step erasure process (see Figure 3):
    1. (a)
      Free expansion of the phase space by a factor 2, leading the system to an undetermined standard state (state S).
    2. (b)
      Quasi-static compression of the phase space by the same factor leading the system from state S to state 0.
  3. 3.
    The first step does not involve any exchanges with the environment, whereas the second dissipates at least $T\ln 2$ of heat, or equivalently at constant internal energy (i.e. at constant temperature for a set of independent data-bits), requires a minimum work. The net balance of the two gives
    $W_{\text{erase/bit}}\geq T\ln 2$ (5)
    which leads, as Brillouin’s law does (Eq.4), to a zero energy balance for Szilard engine. The difference lies in the fact that, according to Landauer, erasure is a necessary step and it is at this very point that the energy cost of data acquisition is paid.
Figure 3: Landauer erasure of a data-bit consisting in a single particle in a two-compartment box. The procedure must be independent of the initial position of the particle: neither a simple shift of the particle towards 0 by means of two pistons keeping constant the volume of the phase space; nor a quasi-static expansion with a single piston is possible. According to Landauer, the first step is necessarily a thermodynamically irreversible free expansion.
Landauer’s law can be examined from different angles.
The first concerns its general validity in the restricted context of data erasure. In reality, even within this context, several points of Landauer’s reasoning are questionable: the imperative to begin erasure with expansion, and the necessity of this expansion to be thermodynamically irreversible. Counterexamples have been provided [39, 40, 41] which invalidate its generality.
The second angle concerns the validity of Landauer’s law on data erasure considered as the missing link between information and energy, something that would replace Brillouin’s law of information. In the following section, we focus on the inconsistencies this idea introduces by confusing information and data.
IV Information versus data
The confusion between information and data is also that between the loss (or deletion) of a bit of information and the erasure of a bit of data. The first refers to the observer’s knowledge, whereas the second only involves a physical object independent of them. A similar confusion in statistical mechanics has already led to inconsistencies. It is therefore not surprising to see others with Landauer’s idea. These inconsistencies make Landauer’s law incompatible with that of Brillouin, that is to say incompatible with the triplet (thermodynamics-statistical mechanics-information theory).
IV.1 Total versus incomplete information
In thermodynamics, setting a data bit into the S state, by removing the partition between two compartments containing a single particle (1st step of Landauer erasure), is reversible or not depending on whether the initial value of the bit is known or not [42] (see Fig. 4). This problem is the same as that of the Gibbs paradoxes and the solution also: thermodynamics considers the case where the observer has an incomplete information about the system, whereas Landauer in his reasoning implicitly assumes that they have the full information at their disposal. There is no solution to this paradox other than those that reintroduce the observer, because in reality, taking these two cases into account (as Bennett does [43], a defender of Landauer’s thesis) already amounts to taking the observer into account.
Figure 4: Setting a data-bit in state S by merging two compartments is thermodynamically irreversible (red path) if the initial value is known ($\Delta S=\ln 2$), but it is thermodynamically reversible (green path) if the initial value is unknown ($\Delta S=0$). This is reminiscent of the Gibbs paradox of mixing. Landauer viewpoint conflicts with thermodynamics.
In addition, the constraint of a single path for erasure, regardless of the initial value of the bit, stems from the requirement to work with unknown values. But this case is thermodynamically reversible. For known values, there is nothing preventing the use of two different reversible paths. The use of a single erasing procedure for known and unknown cases is an arbitrary choice by the observer. Here again lies a clear inconsistency between the intention of theory and its result, the desire to eliminate the observer while reintroducing him.
IV.2 Global concept versus local object
Information is a concept that must not be confuse with its material embodiment. The proof is given by Landauer himself: “Information […] is represented by engraving on a stone tablet, a spin, a charge, a hole in a punched card, a mark on paper, or some other equivalent”[10]. The material embodiment may change but information remains the same. The denial of this difference leads to severe inconsistencies.
A fundamental property of information is that it cannot be given twice. We can duplicate bits of data encoding certain information so that the storage space occupied is twice as large as before, but the corresponding amount of information remains unchanged. Conversely, if multiple copies of the same data bit are erased, the corresponding information will only be lost when the last copy is erased. To assert that information has been erased, one must take into account not only what happened locally, but also the existence or not of a copy somewhere in a more global space.
Information is a global concept, whereas a data-bit is a local object.
It could be argued that this is playing with words and that information can have different meanings. Actually, the common meaning of information as “The imparting of knowledge in general. Knowledge communicated concerning some particular fact, subject, or event” (Oxford Dictionary); is also that of Shannon (if we know the language used for a message, we can compress it further, but indicating the language twice does not allow us to save more space); but this meaning is also that of the defenders of Landauer’s idea (“But what is information? A simple, intuitive answer is “what you don’t already know””[44]. “If someone tells you that the earth is spherical, you surely would not learn much”[45]).
The denial of the difference between information and data and the attempt to materialize information is culminating with the supposed “information-mass equivalence”[46, 12, 47, 48] that leads to paradoxes in relation with the corresponding mass deficit. For instance in the framework of this equivalence, consider a data-bit encoding a bit of information. Duplicate the bit and make the original and the copy physically independent. Erase one of the two. The quantity of information remains unchanged, so that the erased data-bit does not display any mass deficit. Erase the second bit. Information is lost that corresponds to a mass deficit. How does the second bit “knows” that the first has been erased? This contradicts the hypothesis of independence. But if we suppress this hypothesis and assert that independence is impossible, does not that mean that information is not a property of the local data bit, but a property of a global system?
Yet another variation of this paradox. Consider a data-bit, encoding for a bit of information, whose initial value is 0. Erase the data-bit (set its value to 0). There is no longer information stored by the data bit. Erase the bit again. What is the difference between the two erasures? If there is no difference, an eventual mass deficit should not be due to any loss of information. It there is a difference, how the data-bit “knows” that it has been already erased? Only a global observer could.
A data bit has a physical meaning at all scales: from that of a single particle in a two-compartment box, to the macroscopic scale. But information is meaningful only when it involves an observer, information is meaningful only at macroscale. Information, even the smallest piece, is an emergent concept.
IV.3 Pair of states versus path
According to Landauer “Logical irreversibility is associated with physical [thermodynamical] irreversibility”[8]. This is a claim of generality of Landauer’s law and is probably the most problematic point.
Logical irreversibility refers to a loss of information (hence Landauer’s idea of establishing a link with data erasure) we possess about a system. Loss of information that is an increase in the quantity of information we lack to describe the system, or an increase of the uncertainty about it; in other words, an increase in its entropy. Logical irreversibility is simply a positive change in entropy. It is solely related to a change in a state quantity. This is a consequence of the Clausius definition of entropy as a state quantity and of the successive mathematical results of Gibbs and Shannon equating quantity of information and entropy, all condensed in Brillouin’s law of information.
Thermodynamic irreversibility, on the other hand, is a property not linked to a difference between two states, but to the path used to connect them.
How does Landauer arrive at the conclusion that the two irreversibilities are associated? In fact, Landauer forces the path to go through a particular non-differentiable sequence of events, so that a property which initially in the theory is related solely to a peculiar path (thermodynamic irreversibility), becomes also a property of an ordered pair of states (logical irreversibility). This viewpoint either empties the very definition of a state of all meaning, or transforms information and thus entropy into quantities which are not state quantities. In both cases, Landauer’s idea is incompatible with thermodynamics and information theory and leads to inconsistencies.
Consider a system changing from A to B, with $S_{\text{A}}<S_{\text{B}}$. The change is logically irreversible (the information we have decreases), thus (according to Landauer), it is also always thermodynamically irreversible. But the change in the other direction from B to A is logically reversible, thus (according to Landauer), it could be thermodynamically reversible. In summary: the change from A to B would be thermodynamically irreversible, and that from B to A would be thermodynamically reversible. What exactly would mean “thermodynamically reversible” in this context?
In fact, Landauer erasure and Brillouin’s law only lead to the same result in the context of a cyclic data acquisition (such as that of a Szilard machine). Outside of this context, they do not. Brillouin’s law concerns the acquisition of information, Landauer’s law concerns the erasure of data, hence the contradiction if we equate the two. In the context of a Szilard engine and Landauer’s scenario, the data erased before acquisition relates to the engine’s state during the previous cycle. This data is essentially obsolete; it concerns an information about the past. The erased data no longer contains an information about the engine’s current state.
V Conclusive epistemological point
In the literature, Landauer’s law on data erasure is generally presented as a decisive improvement to information theory that provides us with the missing key to understanding the link between information and thermodynamics. For instance, one can read: “The Landauer principle is one of the cornerstone of the modern theory of information” (Herrera [12]).
This situation is very strange because there is absolutely no experiment, no observation, that can be explained by Landauer’s law and that would not be explained without it. On one side, the link between information and energy, as established by Maxwell or Szilard engines, is perfectly explained by Brillouin’s law. On the other side, the erasure of data bits, whatever their form, is also perfectly explained by thermodynamics. Landauer’s law adds absolutely nothing to the theory, except inconsistencies that make it incompatible with thermodynamics and information theory. So, how do we conceive its popularity? It is as if an explanation that consists solely of deducing observations (past and future) from a set of theoretical statements (principles and laws) is insufficient. Yet, that is precisely the only thing expected of theory.
The popularity of Landauer’s law can only be due to a misconception of what are concepts in physics. There is no physical theory without concepts. But concepts such as energy and information (but also that of field and many others) are not material entities, although they are fully physical, in the sense that they play a crucial role in current physical theories. They are, in fact, conventions. The meaning of any concept, what it encompasses, its definition, is a convention because we cannot see it and consequently, establish a direct link between what we see and we mean. We cannot show a concept to someone.
“Information is physical” (Landauer [9]). “Information is not a disembodied abstract entity” (Landauer [10]). Landauer’s law on data erasure is in reality an attempt to materialize information, which, without Landauer, would remain an abstract concept. According to his idea, the smallest piece of information is actually localized and materialized under the form of a data-bit. In this sense, Landauer’s idea belongs to the reductionist branch of materialism. It is in direct continuity with the initial intention of Gibbs’ statistical mechanics. Here, the purpose is neither to contest materialism, nor to deny the utility of reductionism when it allows an economy of thought. It is simply to deny reductionism as a profession of faith or a creed.
The concept of energy is widely used in physics although it is a pure concept. There is no such thing as an “energy particle” (an object) that could exist independently of us. Energy is like a universal. And nobody has a problem with that. Why not consider information in the same way?
Attached to materialism and to “the scientific conception of the world”[13] (neo-positivism) is the idea that the only two sources of knowledge are that underlying mathematics, with no other synthetic a priori knowledge, and that of observations. The former imposes a strict requirement of consistency in the statements of the theory, while the latter obliges us to take into account the role of the observer, since there is no observation without observer.
This is where a potential problem of misunderstanding lies. Accounting for the observer introduces information. A concept which is emergent since the observer is only meaningful at the macroscopic scale. Hence the danger of falling into a form of holism that would consist of considering emergence as synthetic a priori knowledge. In other words, precisely what neo-positivism intends to reject. This apparent contradiction introduced by emergence into materialism probably explains the search for a reductionist solution like Landauer’s.
Actually, a theoretical statement involving an emergent property, such as the principle of maximum entropy, is not a synthetic a priori knowledge, it is a convention. A convention that is exactly of the same nature as that of the first principle of thermodynamics or Euclidian geometry [18, 16]. The choice of one convention or another is guided solely by its convenience and in particular by the economy of thought [14] it provides: the simplest is the best. The prior probability distribution of phases of a system, which is directly deduced from the principle of maximum entropy, is in this sense also a convention (it is directly derived from a convention). It does not pretend to reflect the reality of the system. It is not the system that maximizes the probability distribution of its phases, but our conventional representation of the system that must do so if we want to be rational.
Emergentism is generally and historically divided into two categories [49]: ontological and epistemological. The former considers it is inherent in the essence of things, the latter that it is only imposed to us by our necessary limited knowledge. However, with regard to experience, both lead to the same result. The difference is a metaphysical question out of the scope of science. The neo-positivist position [13, 16], inspired by Poincaré’s conventionalism [18], which is defended here, makes it possible to avoid this problem and constitutes a third way which could be called “conventional emergentism” and is not so far removed from pragmatism [50, 51].